OpenStudy (mathmath333):

A bouquet has to be formed from 18 different flowers so that it should contain not less than 3 flowers. How many ways are there of doing this in ?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{ A bouquet has to be formed from 18 different flowers so}\ \hspace{.33em}\\~\\ & \normalsize \text{ that it should contain not less than 3 flowers. How many}\ \hspace{.33em}\\~\\ & \normalsize \text{ ways are there of doing this in ?}\ \hspace{.33em}\\~\\ & a.)\ 524288 \hspace{.33em}\\~\\ & b.)\ 262144 \hspace{.33em}\\~\\ & c.)\ 261972 \hspace{.33em}\\~\\ & d.)\ \normalsize \text{none of these} \hspace{.33em}\\~\\ \end{align}}\)

Parth (parthkohli):

I hope this isn't a circular permutations question, but just a normal combinations wala question.

OpenStudy (mathmath333):

Is it \(\large \color{black}{\begin{align} & 2^{18}-(1+2^{1}+2^{2}+2^{3})\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

is answer option c.)

ganeshie8 (ganeshie8):

The problem is equivalent to forming subsets with cardinality > 2 from a set with cardinality 18

Parth (parthkohli):

\[2^{18} - \left(\binom{18}{0} + \binom{18}{1}+\binom{18}{2}\right)\]

Parth (parthkohli):

oh no he beat me to it

Parth (parthkohli):

one day I'll get my internet speed upgraded... one day...

OpenStudy (mathmath333):

hey but the answer given is option d.) none

ganeshie8 (ganeshie8):

lol PK I heard ACT provides higher speeds; mine is airtel, though http://www.acttv.in/index.php/products/act-broadband

ganeshie8 (ganeshie8):

@mathmath333 does it give any explanation ? we might get a different answer if we interpret it differently

OpenStudy (mathmath333):

wait

ganeshie8 (ganeshie8):

Anyways, this is a trivial question. I feel not worth spending time... must be a typo or a mistake of author..

OpenStudy (mathmath333):

It doesn't has any explanation \(\large \color{black}{\begin{align} & 2^{18}-(\dbinom{18}{0}+\dbinom{18}{1}+\dbinom{18}{2}+\dbinom{18}{3})=261158\hspace{.33em}\\~\\ \end{align}}\)

ganeshie8 (ganeshie8):

you can have 3 flowers, so you should not subtract that

Parth (parthkohli):

I know about ACT but I'm from Delhi. And I have Airtel too.

OpenStudy (mathmath333):

ok i assume it to be typo

Parth (parthkohli):

OH YOU'RE FROM BANGALORE. I'm so jealous, Ganeshie.

imqwerty (imqwerty):

yea c is correct ok the number of ways wuld be 18C0+18C1+18C2 + 18C3+18C4+18C5........18C18-18C0 - 18C1 -18C2 \[2^\left( 18 \right)-1-18-153\] =261972